Balow are h ieney numbers of l elcted fiom a o ai tan, Find the range vaniance, and standard deviation for the given sample data What do the results tell us? 41 63 38 24 53 73 1 90 33 95 26 Range(Round to one decimal place as needed)
Answer #1

answer is Range=94

here the minimum =1 and maximum=95

the range=maximum-minimum=95-1=94

variance=sum((x-mean)2/(n-1))=8483.64/10=848.4 ( one decimal approximation)

standard deviation=sqrt(variance)=sqrt(848.4)=29.1

following information has been generated for answering above

sn x x-mean (x-mean)2
1 1 -47.82 2286.7524
2 24 -24.82 616.0324
3 26 -22.82 520.7524
4 33 -15.82 250.2724
5 38 -10.82 117.0724
6 41 -7.82 61.1524
7 53 4.18 17.4724
8 63 14.18 201.0724
9 73 24.18 584.6724
10 90 41.18 1695.7924
11 95 46.18 2132.5924
sum= 537 -0.02 8483.6364
n= 11.0
mean=sum/n= 537/11=48.82